# Designing Consensus-Based Distributed Filtering over Directed Graphs

### v1

@@ 3.2 Two Estimation Errors
    The measurement estimation error $\varepsilon^{(l)}_{k}$ is given by
    $\varepsilon^{(l)}_{k}=G^{l}\varepsilon^{(0)}_{k},$ (12)
    where
    $G$ $=I_{Nr}-\Lambda(\mathcal{L}\otimes I_{r}+B),~{}~{}~{}~{}~{}$ (13)
    $\Lambda=\text{diag}\{\Lambda_{1},\ldots,\Lambda_{N}\},{}{}{}{}{}{}{}{}{}{}{}{}{}$ (14)
    $B=\text{diag}\{B_{1},\ldots,B_{N}\},{}{}{}{}{}{}{}{}{}{}{}{}{}$ (15)
    and
    $\varepsilon^{(0)}_{k}$ $=(I_{N}\otimes CA)e_{k-1}-1_{N}\otimes\nu_{k}$ (16)
@@
    The expectation value $E\{\varepsilon^{(l)}_{k}\}$ and $P_{\varepsilon,k}$ are given by
    $E\{\varepsilon^{(l)}_{k}\}=G^{l}(I_{N}\otimes CA)E\{e_{k-1}\},$ (17)
    and
@@
    $~{}~{}~{}~{}+G^{l}(I_{N}\otimes C)(U_{N}\otimes Q)(I_{N}\otimes C)^{T}(G^{l})^{T}$
    $~{}~{}~{}~{}+G^{l}(U_{N}\otimes R)(G^{l})^{T},$
    respectively.
The proof of Proposition 1 is given in Appendix A.
Lemma 5.
Under the assumption that $\rho(G)<1$ , as $l$ tends to infinity, it holds
$\lim_{l\to\infty}E\{\varepsilon^{(l)}_{k}\}=0,$ (19)
and
$\lim_{l\to\infty}P_{\varepsilon,k}=0.{}{}{}$ (20)
The proof of Lemma 5 is given in Appendix B.
Remark 5.
Proposition 1 shows that $\varepsilon^{(l)}_{k}$ is a random variable. If $e_{k-1}$, $\omega_{k-1}$, and $\nu_{k}$ are Gaussian, $\varepsilon^{(l)}_{k}$ is Gaussian with the mean in (17) and the covariance $P_{\varepsilon,k}$ in (18). Lemma 5 demonstrates that as $l$ tends to infinity, $\varepsilon^{(l)}_{k}$ converges to 00. The state estimation error $e_{k}$ and the measurement estimation error $\varepsilon^{(l)}_{k}$ interact and influence each other.
Proposition 2.
The estimation error $e_{k}$ and the estimation error covariance $P_{k|k}=E\{e_{k}e^{T}_{k}\}$ are
$e_{k}=\mathcal{A}(l)e_{k-1}-\mathcal{B}(l)(1_{N}\otimes\omega_{k-1})+\mathcal{D}(l)(1_{N}\otimes\nu_{k}),$ (21)
and
$P_{k|k}$ $=\mathcal{A}(l)P_{k-1|k-1}\mathcal{A}^{T}(l)+\Phi(l),$ (22)
respectively, where
$\Phi(l)=\mathcal{B}(l)(U_{N}\otimes Q)\mathcal{B}^{T}(l)+\mathcal{D}(l)(U_{N}\otimes R)\mathcal{D}^{T}(l),$ (23)
$\mathcal{A}(l)=I_{N}\otimes(A-KCA)+(I_{N}\otimes K)G^{l}(I_{N}\otimes CA),$ (24)
@@
The proof of Proposition 2 is given in Appendix C.
Remark 6.
Based on Proposition 2, two parameters still need to be designed: one is the parameter $\mu_{ij}$ in $G$, and the other is the fusion step $l$. Observing (24), $\rho(G)<1$ ensures the convergence of the filter. Subsequently, we focus on the design of the parameters $\mu_{ij}$ and $l$.
3.3 Parameter Design for $\mu_{i}$
The objective of designing the parameter $\mu_{i}$ is to ensure that $\rho(G)<1$, and it can guarantee the convergence of the measurement estimate error and the state estimate error. Next, two methods are provided for the design of the parameter $\mu_{i}$.
3.3.1 Distributed Design
A distributed design approach for $\mu_{ij}$ is proposed to circumvent the utilization of the global topology information for each sensor. First, the elements of $G$ are presented, laying the foundation for subsequent design considerations.
Define the submatrix $G_{[ij]}$ as
$G_{[ij]}$ $=[0_{r\times r(i-1)},I_{r},0_{r\times r(N-i)}]G$
×$~{}~{}~{}~{}\times[0^{T}_{r\times r(j-1)},I^{T}_{r},0^{T}_{r\times r(N-j)}]^{T},$
where $[ij]$ represents the $i$-th row and $j$-th column index of the corresponding matrix block. The detailed submatrices are
$G_{[ii]}$ $=\text{diag}\{g_{ii,1},\dots,g_{ii,N}\},$
and
$G_{[ij]}$ $=\text{diag}\{g_{ij,1},\dots,g_{ij,N}\},$
@@ 3.3.1 Distributed Design
$~{}~{}~{}~{}~{}g_{[ii,l]}$ $=(1-\mu_{il}(l_{ii}+a_{il}))\otimes I_{r_{l}},$ (27)
and
$g_{[ij,l]}$ $=(-\mu_{il}l_{ij})\otimes I_{r_{l}}.~{}~{}~{}~{}~{}~{}~{}$ (28)
Consider the row sum of the block matrix, and it holds
$\sum^{n}_{j=1}g_{[ij,l]}$ $=\Big{(}1-\mu_{il}\sum^{N}_{j=1}{l_{ij}}-\mu_{il}a_{il}\Big{)}\otimes I_{r_{l}}$ (29)
$=\Big{(}1-\mu_{il}a_{il}\Big{)}\otimes I_{r_{l}}.$
Next, a reasonable range of the parameter $\mu_{ij}$ is provided to make $\rho(G)<1$.
Lemma 6.
Under Assumption 1, if $\mu_{il}$ satisfies
$0<\mu_{il}\leq\frac{1}{l_{ii}+a_{il}},$
then it holds $\rho(G)<1$.
The proof of Lemma 6 is given in Appendix D.
Remark 7.
The distributed design method can be employed to design parameter $\mu_{ij}$ by utilizing the local information instead of relying on global topology information [10, 3]. Additionally, It is suggested to design $\mu_{ij}$ such that $1-\mu_{il}(l_{ii}+a_{il})$ approaches 00, leading to a smaller $\rho(G)$ with a higher probability.
3.3.2 Unified Design
A unified design method is also provided for undirected graphs. Consider the scenario where all $\mu_{i}$ are identical, denoted as $\mu$. Subsequently, the symbol $G$ is redefined as follows:
$G$ $=I_{Nr}-\mu(\mathcal{L}\otimes I_{r}+B).$
@@ 3.3.2 Unified Design
Based on Lemma 1, $G$ is positive definite. By selecting $\mu<\frac{1}{\rho(\mathcal{L}\otimes I_{r}+B)}$, it is ensured that $\rho(G)<1$. This method needs the global communication topology information.
3.4 Parameter Design for $l$
@@
This subsection aims to design the parameter $l$ to ensure the convergence of the distributed estimators. By employing Lemma 2 and considering (24), it is evident that $\rho(\mathcal{A}(l))<1$ can guarantee the convergence of the distributed filter. The following results provide a lower bound of the consensus step $l$.
Theorem 1.
@@
Under Assumption 1 and 2, the estimation error covariance $P_{i,k|k}$ is uniformly upper-bounded, if $\|G\|_{2}<1$ and $l>l_{0}$, where
$l_{0}=\text{log}_{\|G\|_{2}}\frac{1-\|A-KCA\|_{2}}{\|K\|_{2}\|CA\|_{2}}.$ (30)
The proof of Theorem 1 is given in Appendix E.
Remark 8.
@@
For Theorem 1, there are two elements that decide the value of $l_{0}$, i.e., $\|G\|_{2}$ and $\bar{K}=\frac{1-\|A-KCA\|_{2}}{\|K\|_{2}\|CA\|_{2}}$. On one hand, the term $\|G\|_{2}$ is relevant to the communication graph and the parameter $\mu_{i}$. The smaller $\|G\|_{2}$, the smaller $l_{0}$. On the other hand, the term $\bar{K}$ indicates the stability margin of the estimator. The larger $\bar{K}$, the smaller $l_{0}$.
4 Performance Analysis
@@
In this section, the centralized estimator is first introduced. Then, the steady-state performance gap between the centralized filter and the proposed distributed filter is analyzed with the increasing fusion step $l$. Additionally, the transient performance is also taken into consideration.
4.1 Centralized Estimator
Consider a centralized estimator as follows:
$\hat{x}_{c,k|k-1}=A\hat{x}_{c,k-1|k-1},$ (31)
$\hat{x}_{c,k|k}=\hat{x}_{c,k|k-1}+K(y_{k}-C\hat{x}_{c,k|k-1}),$ (32)
where $y_{k}$, $K$, and $C$ are given in (6).
For the centralized estimator, denote the estimation error as $e_{c,k|k}=\hat{x}_{c,k|k}-x_{k}$, the corresponding estimation error covariance as $P_{c,k|k}=E\{e_{c,k|k}e^{T}_{c,k|k}\}$, the augmented estimation error as $e_{cc,k}=1_{N}\otimes e_{c,k}$, and the augmented estimation error covariance as $P_{cc,k|k}=E\{e_{cc,k|k}e^{T}_{cc,k|k}\}=U_{N}\otimes P_{c,k|k}.$
@@ 4.1 Centralized Estimator
By combining the estimator (31), (32) and the dynamical system (1), the estimation error $e_{c,k}$ is
$e_{c,k}$ $=(A-KCA)\hat{x}_{c,k-1|k-1}+Ky_{k}-Ax_{k-1}-\omega_{k-1}$ (33)
$=(A-KCA)e_{c,k-1}-(I-KC)\omega_{k-1}+K\nu_{k}.$
Then, $P_{c,k|k}=E\{e_{c,k|k}e^{T}_{c,k|k}\}$ can be computed as
$P_{c,k|k}$ $=(A-KCA)P_{c,k-1|k-1}(A-KCA)^{T}$ (34)
$~{}~{}~{}+(I-KC)Q(I-KC)^{T}+KRK^{T}.$
Similarly, the augmented estimation error of the centralized estimator $e_{cc,k}=1_{N}\otimes e_{c,k}$ and the corresponding estimation error covariance $P_{cc,k|k}$ can be obtained as
$e_{cc,k}=\mathcal{A}_{cc}e_{cc,k-1}-\mathcal{B}_{cc}(1_{N}\otimes\omega_{k-1})+\mathcal{D}_{cc}(1_{N}\otimes\nu_{k}),$ (35)
and
$P_{cc,k|k}$ $=\mathcal{A}_{cc}P_{cc,k-1|k-1}\mathcal{A}^{T}_{cc}+\Phi_{cc},~{}~{}~{}~{}~{}~{}~{}~{}$ (36)
respectively, where
$\Phi_{cc}=\mathcal{B}_{cc}(U_{N}\otimes Q)\mathcal{B}^{T}_{cc}+\mathcal{D}_{cc}(U_{N}\otimes R)\mathcal{D}^{T}_{cc},$ (37)
@@
Remark 9.
This centralized estimator utilizes the fixed gain matrix $K$, resulting in a suboptimal filter at every step. Nevertheless, as the time step $k$ approaches infinity, $P_{c,k|k}$ converges to $P$ in (7). Therefore, the relations between the steady-state performance of $P_{i,k|k}$ and $P_{c,k|k}$ reflect those between the steady-state performance of $P_{i,k|k}$ and $P$.
4.2 Convergence Analysis
This section investigates the steady-state performance gap between $P_{cc,k|k}$ and $P_{k|k}$, and sheds light on the influence of the fusion step $l$ on the performance. Section 3.3 has shown that the parameter $\mu_{ij}$ can be designed to ensure that $\rho(G)<1$. Therefore, we make the following assumption.
Assumption 3.
@@ 4.2 Convergence Analysis
Theorem 2.
Under Assumption 1, 2, and 3, $P_{k|k}$ and $P_{cc,k|k}$ converge to the unique solutions of the discrete-time Lyapunov equations (DLE)
$P^{(l)}$ $=\mathcal{A}(l)P^{(l)}\mathcal{A}^{T}(l)+\Phi(l),$ (41)
@@
The proof of Theorem 2 is given in Appendix F.
To evaluate the effect of the fusion step $l$ on the steady-state performance of the proposed filter, we introduce the following notations
$\bar{\mathcal{A}}(l)=\mathcal{A}(l)-\mathcal{A}_{cc}=(I_{N}\otimes K)G^{l}(I_{N}\otimes CA),$
@@
The proof of Lemma 7 is given in Appendix G.
Theorem 3.
Under Assumption 1 and 2, there exist constants $M_{1},M_{2}>0$, such that
@@
The proof of Theorem 3 is given in Appendix H.
Remark 10.
Two conditions related to the communication graph are presented, since the mere satisfaction of a spectral radius of the matrix less than 1 does not ensure the spectral norm of the matrix less than 1. For an undirected graph, the spectral radius $\rho(G)$ is equal to the spectral norm $\|G\|_{2}$, and the conditions of item (1) and (2) in Theorem 3 are identical. For both items, as the fusion step $l$ tends to infinity, $\|P^{(l)}-P_{cc}\|_{2}$ converges to 00 according to (45) and (46).
4.3 Transient Performance
The transient performance of the proposed filter is analyzed in the following.
Lemma 8.
Under Assumption 3, as $l$ tends to infinity, it holds that
$\lim_{l\to\infty}\mathcal{A}(l)=\mathcal{A}_{cc},$
$\lim_{l\to\infty}\mathcal{B}(l)=\mathcal{B}_{cc},$
and
$\lim_{l\to\infty}\mathcal{D}(l)=\mathcal{D}_{cc}.$
The proof of Lemma 8 is given in Appendix I.
Theorem 4.
Under Assumption 1, 2, and 3, and considering the case that $P_{k-1|k-1}=P_{cc,k-1|k-1}$, one has
$\lim_{l\to+\infty}P_{i,k|k}=P_{c,k|k}.$ (47)
The proof of Theorem 4 is given in Appendix J.
Remark 11.
Based on the structural characteristics of $\mathcal{A}(l)$, $\mathcal{B}(l)$, and $\mathcal{D}(l)$, Lemma 8 shows that the matrices $\mathcal{A}(l)$, $\mathcal{B}(l)$, and $\mathcal{D}(l)$ approach to the centralized matrices $\mathcal{A}_{cc}$, $\mathcal{B}_{cc}$, and $\mathcal{D}_{cc}$, respectively, as the fusion step tends to infinity. Then, it is demonstrated that the transient performance exhibits similar results. In other words, as the fusion step increases, the performance of the proposed filter approaches that of the centralized filter. By considering both the transient properties and the convergence properties, we can gain comprehensive insights into the performance of the proposed algorithm.
5 Simulations

### v2

@@ III-B Two Estimation Errors
    The measurement estimation error $\varepsilon^{(l)}_{k}$ is
    $\varepsilon^{(l)}_{k}=G^{l}\varepsilon^{(0)}_{k}$,
    where $G=I_{Nr}-\Lambda(\mathcal{L}\otimes I_{r}+B)$, $G^{l}$ represents the $l$-th power of $G$, $\Lambda=\text{diag}\{\Lambda_{1},\ldots,\Lambda_{N}\}$, $B=\text{diag}\{B_{1},\ldots,B_{N}\}$, and
    $\varepsilon^{(0)}_{k}$ $=(I_{N}\otimes CA)e_{k-1}-1_{N}\otimes\nu_{k}$
@@
    The measurement estimation error covariance $P_{\varepsilon,k}$ is
@@
    $~~~~+G^{l}(I_{N}\otimes C)(U_{N}\otimes Q)(I_{N}\otimes C)^{T}(G^{l})^{T}$~~
    $~~~~+G^{l}(U_{N}\otimes R)(G^{l})^{T}.$
  3. 3.
    The state estimation error $e_{k}$ is
    $~~e_{k}=\mathcal{A}(l)e_{k-1}-\mathcal{B}(l)(1_{N}\otimes\omega_{k-1})+\mathcal{D}(l)(1_{N}\otimes\nu_{k}),$
    where
    $~\mathcal{A}(l)=I_{N}\otimes(A-KCA)+(I_{N}\otimes K)G^{l}(I_{N}\otimes CA),$ (8)
@@
  4. 4.
    The state estimation error covariance $P_{k|k}$ is
    $~~P_{k|k}=\mathcal{A}(l)P_{k-1|k-1}\mathcal{A}^{T}(l)+\Phi(l),$ (11)
@@
    $\Phi(l)=\mathcal{B}(l)(U_{N}\otimes Q)\mathcal{B}^{T}(l)+\mathcal{D}(l)(U_{N}\otimes R)\mathcal{D}^{T}(l).$ (12)
The proof of Proposition 1 is given in Appendix A.
Remark 4
Proposition 1 shows that $\varepsilon^{(l)}_{k}$ is a random variable with the mean $E\{\varepsilon^{(l)}_{k}\}=G^{l}(I_{N}\otimes CA)E\{e_{k-1}\}$ and the covariance $P_{\varepsilon,k}$. Under the assumption that $\rho(G)<1$, as $l$ tends to infinity, it holds $\lim_{l\to\infty}E\{\varepsilon^{(l)}_{k}\}=0$ and $\lim_{l\to\infty}P_{\varepsilon,k}=0$. The state estimation error $e_{k}$ and the measurement estimation error $\varepsilon^{(l)}_{k}$ interact and influence each other. To improve the filter performance, two parameters will be designed: one is the parameter $\mu_{ij}$ in $G$, and the other is the fusion step $l$.
III-C Parameter Design for $\mu_{i}$
This subsection will design the parameter $\mu_{i}$ to ensure the convergence of two estimation errors.
III-C1 Distributed Design
We propose a distributed design approach for $\mu_{ij}$ to circumvent the utilization of the global topology information for each sensor. First, define the submatrix $G_{[ij]}$ as $G_{[ij]}=[0_{r\times r(i-1)},I_{r},0_{r\times r(N-i)}]G[0^{T}_{r\times r(j-1)},I^{T}_{r},0^{T}_{r\times r(N-j)}]^{T},$ where $[ij]$ represents the $i$-th row and $j$-th column index of the corresponding matrix block. The detailed submatrices are $G_{[ii]}=\text{diag}\{g_{ii,1},\dots,g_{ii,N}\}$ and $G_{[ij]}=\text{diag}\{g_{ij,1},\dots,g_{ij,N}\}$, where
$~~~~g_{[ii,l]}$ $=(1-\mu_{il}(l_{ii}+a_{il}))\otimes I_{r_{l}},$ (13)
$g_{[ij,l]}$ $=(-\mu_{il}l_{ij})\otimes I_{r_{l}},$~~~~~~~ (14)
and $l_{ij}$ is the $i$-th row and $j$-th column element of the Laplacian matrix $\mathcal{L}$. Next, we introduce a method to select the parameter $\mu_{ij}$ from a reasonable range, given by:
$0<\mu_{il}\leq\frac{1}{l_{ii}+a_{il}}.$~~~~~~~~~~~~ (15)
We provide the following lemma to illustrate the choice of $\mu_{il}$ satisfying (15) can ensure $\rho(G)<1$.
Lemma 5
Under Assumption 2, if $\mu_{il}$ satisfies (15) then it holds $\rho(G)<1$.
The proof of Lemma 5 is given in Appendix B.
III-C2 Unified Design
@@
A unified design method is also provided for undirected graphs. Consider the scenario where all $\mu_{i}$ are identical, denoted as $\mu$. The symbol $G$ is then redefined as $G=I_{Nr}-\mu(\mathcal{L}\otimes I_{r}+B).$ Based on Lemma 1, $G$ is positive definite. By selecting $\mu<\frac{1}{\rho(\mathcal{L}\otimes I_{r}+B)}$, it is ensured that $\rho(G)<1$. This method requires the global communication topology information.
III-D Parameter Design for $l$
@@
This subsection aims to provide the lower bound of the fusion step number $l$ to ensure the convergence of the distributed estimators.
Theorem 1
Under Assumptions 1 and 2, if $\|G\|_{2}<1$ and $l>l_{0}$, where
$~l_{0}=\text{log}_{\|G\|_{2}}\frac{1-\|A-KCA\|_{2}}{\|K\|_{2}\|CA\|_{2}},$ (16)
@@
then $\mathcal{A}(l)$ is Schur stable, and the estimation error covariance $P_{i,k|k}$ is uniformly upper-bounded.
The proof of Theorem 1 is given in Appendix C.
Remark 5
@@
From Theorem 1, the value $l_{0}$ is determined by two factors: $\|G\|_{2}$ and $\bar{K}=\frac{1-\|A-KCA\|_{2}}{\|K\|_{2}\|CA\|_{2}}$. On one hand, the term $\|G\|_{2}$ is related to the communication graph and the parameter $\mu_{i}$. The smaller $\|G\|_{2}$ is , the smaller $l_{0}$ becomes. On the other hand, the term $\bar{K}$ indicates the stability margin of the estimator. The larger $\bar{K}$ is, the smaller $l_{0}$ becomes.
IV Performance Analysis
@@
This section analyzes the performance of the proposed distributed filter. Specifically, we present the steady-state performance gap between the centralized filter and the proposed distributed filter as the fusion step number $l$ increases.
IV-A Centralized Estimator
Consider a centralized estimator for comparison, given by:
$\hat{x}_{c,k|k}=A\hat{x}_{c,k-1|k-1}+K(y_{k}-CA\hat{x}_{c,k-1|k-1}),$ (17)
@@ IV-A Centralized Estimator
Denote the estimation error as $e_{c,k|k}=\hat{x}_{c,k|k}-x_{k}$, the estimation error covariance as $P_{c,k|k}=E\{e_{c,k|k}e^{T}_{c,k|k}\}$, the augmented estimation error as $e_{cc,k}=1_{N}\otimes e_{c,k}$, and the augmented estimation error covariance as $P_{cc,k|k}=E\{e_{cc,k|k}e^{T}_{cc,k|k}\}=U_{N}\otimes P_{c,k|k}.$ By combining the estimator (17) and the dynamical system (1), $e_{cc,k}=1_{N}\otimes e_{c,k}$ can be computed as
$~~~e_{cc,k}=\mathcal{A}_{cc}e_{cc,k-1}-\mathcal{B}_{cc}(1_{N}\otimes\omega_{k-1})+\mathcal{D}_{cc}(1_{N}\otimes\nu_{k}),$ (18)
where
@@
The corresponding estimation error covariance $P_{cc,k|k}$ is
$P_{cc,k|k}$ $=\mathcal{A}_{cc}P_{cc,k-1|k-1}\mathcal{A}^{T}_{cc}+\Phi_{cc},$~~~~~~~~~~ (22)
where
$\Phi_{cc}=\mathcal{B}_{cc}(U_{N}\otimes Q)\mathcal{B}^{T}_{cc}+\mathcal{D}_{cc}(U_{N}\otimes R)\mathcal{D}^{T}_{cc}.$ (23)
Since we analyze the steady-state performance gap, the centralized filter with the fixed gain will converge to the optimal solution.
IV-B Convergence Analysis
This section investigates the steady-state performance gap between $P_{cc,k|k}$ and $P_{k|k}$, and sheds light on the influence of the fusion step number $l$ on the performance. Section III-C has shown that the parameter $\mu_{ij}$ can be designed to ensure that $\rho(G)<1$. Therefore, we make the following assumption.
Assumption 3
@@ IV-B Convergence Analysis
Remark 6
Assumption 3 can be satisfied through the proposed parameter design method, as established in Lemma 5 and Theorem 1. In other words, this assumption indicates that the parameters have been appropriately designed.
Theorem 2
Under Assumptions 1, 2, and 3, $P_{k|k}$ and $P_{cc,k|k}$ converge to the unique solutions of the discrete-time Lyapunov equations (DLE)
$P^{(l)}$ $=\mathcal{A}(l)P^{(l)}\mathcal{A}^{T}(l)+\Phi(l),$~~ (24)
@@
The proof of Theorem 2 is given in Appendix D.
To evaluate the effect of the fusion step number $l$ on the steady-state performance, we introduce the following notations
$\bar{\mathcal{A}}(l)=\mathcal{A}(l)-\mathcal{A}_{cc}=(I_{N}\otimes K)G^{l}(I_{N}\otimes CA),$ (26)
@@
The proof of Lemma 6 is given in Appendix E.
Theorem 3
Under Assumptions 1 and 2, there exist constants $M_{1},M_{2}>0$, such that
@@
The proof of Theorem 3 is given in Appendix F.
Remark 7
Two conditions related to the communication graph are presented, since ensuring that the spectral radius of the matrix is less than 1 does not necessarily imply that its spectral norm is also less than 1. For an undirected graph, the spectral radius $\rho(G)$ equals the spectral norm $\|G\|_{2}$, and the conditions in item (1) and (2) of Theorem 3 are identical. In both cases, as the fusion step number $l$ approaches infinity, $\|P^{(l)}-P_{cc}\|_{2}$ converges to 0, as demonstrated in (29) and (30).
V Simulations